# Scenario analysis with multivariate Bayesian machine learning models

### v1

@@ 2.1. Multivariate system estimation
To estimate the multivariate (reduced form) model in Equation (1) efficiently from a computational perspective, we rely on a conditional representation of its $n$ equations. Let $\bm{e}_{i}$ of size $1\times n$ denote the $i$th row of an identity matrix $\bm{I}_{n}$, and $\bm{E}_{i}$ of size $(n-1)\times n$ is the matrix resulting from deleting the $i$th row of $\bm{I}_{n}$. Using $\bm{y}_{-it}=\bm{E}_{i}\bm{y}_{t}$ we may write $\bm{y}_{t}=\bm{e}_{i}^{\prime}y_{it}+\bm{E}_{i}^{\prime}\bm{y}_{-it}$. Under the assumptions of Equation (1) one may derive the conditional distribution: [^1] Earlier related contributions often either use a mapping between the structural and reduced form of the VAR to enable equation-by-equation estimation (see, e.g., Hauzenberger et al., 2024c), or rely on factor models for the reduced form errors (see, e.g., Clark et al., 2023). These approaches come with computational and inferential advantages and disadvantages. The former is simple to implement but requires parameterizing the structural form which may cause associated issues such as breaking order-invariance of the equations. The latter allows for order-invariant inference but gives rise to the usual identification challenges of factor models. Our proposed approach allows for directly working with the reduced form model, and obtaining order-invariant inference.
@@
which is a Gaussian with variance $\varsigma_{it}^{2}=(\bm{e}_{i}\bm{\Sigma}_{t}^{-1}\bm{e}_{i}^{\prime})^{-1}$ and mean ${\mu}_{it}=\varsigma_{it}^{2}(\bm{e}_{i}\bm{\Sigma}_{t}^{-1}(\bm{F}(\bm{x}_{t})-\bm{E}_{i}^{\prime}\bm{y}_{-it}))$. Indeed, this distribution is equivalent to a more common representation of the conditional multivariate Gaussian (see, e.g., Cong et al., 2017, Section 2, Equation 5). The mean can alternatively be written as ${\mu}_{it}=f_{i}(\bm{x}_{t})+\varsigma_{it}^{2}(\bm{e}_{i}\bm{\Sigma}_{t}^{-1}\bm{E}_{i}^{\prime})(\bm{y}_{-it}-\bm{E}_{i}\bm{F}(\bm{x}_{t}))$, and we define $\tilde{\mu}_{it}=\varsigma_{it}^{2}(\bm{e}_{i}\bm{\Sigma}_{t}^{-1}\bm{E}_{i}^{\prime})(\bm{y}_{-it}-\bm{E}_{i}\bm{F}(\bm{x}_{t}))$, i.e., $\mu_{it}=f_{i}(\bm{x}_{t})+\tilde{\mu}_{it}$.
@@
which can be used in a Gibbs sampler to update the conditional mean relationships by looping through equations $i=1,\ldots,n$. This approach is similar to the one of Esser et al. (2024) and allows to treat each equation of the multivariate system individually, conditional on all other equations.
@@ 2.2. Bayesian Additive Regression Trees
The approach we discuss in Section 3 works with any implementation of multivariate models with jointly Gaussian errors. That is, assuming a linear functional form for $\bm{F}(\bm{x}_{t})$ results in a standard BVAR. [^2] When we consider linear versions of our model for comparisons, we implement this setting with $\bm{F}(\bm{x}_{t})=\bm{A}\bm{x}_{t}$ where $\bm{A}$ is an $n\times k$ matrix of reduced form VAR coefficients. We assume a horseshoe prior with a single global shrinkage parameter on these coefficients, see also Hauzenberger et al. (2024b). Equation (2) can be used to update the VAR coefficients equation-by-equation from their usual Gaussian posteriors. In case we choose to treat $\bm{F}(\bm{x}_{t})$ nonparametrically, several options are available. Due to its versatility and the established favorable empirical properties we mentioned earlier, we use BART to approximate the equation-specific functions in our applied work. That is, we consider a sum of $s=1,\ldots,S,$ tree functions $\ell_{is}(\bm{x}_{t}\>|\>\mathcal{T}_{is},\bm{\mathrm{m}}_{is})$:
@@ 2.4. Posterior distributions and sampling algorithm
on a $t$-by-$t$ basis. The posterior distribution of the outlier probability is $\mathfrak{p}\>|\>\bullet\sim\mathcal{B}(a_{\mathfrak{p}}+T_{\text{o}},b_{\mathfrak{p}}+T-T_{\text{o}})$, with the total number of observations classified as outliers denoted by $T_{\text{o}}=\sum_{t=1}^{T}\mathbb{I}(s_{t}\neq 1)$, where $\mathbb{I}(\bullet)$ is an indicator function that yields $1$ if its argument is true and 00 otherwise.
This completes our modeling framework and algorithmic implementation. We turn to the issue of predictive and structural inference next.
3. Predictive simulation
Define a vector $\bm{\Xi}$ that contains all coefficients and latent variables necessary to parameterize Equation (1). At time $\tau$, the one-step-ahead predictive distribution is given by the integral:
[…]
where $\mathcal{I}$ denotes the respective information set used to infer $\bm{\Xi}$. For out-of-sample forecasts, $\mathcal{I}$ is typically given by $\{\bm{y}_{t}\}_{t=1}^{\tau}$, and we are interested in predicting $\bm{y}_{\tau+h}\>|\>\{\bm{y}_{t}\}_{t=1}^{\tau}$ for $h=1,2,\ldots,$ steps ahead. In other cases we may also want to condition on the full information set, i.e., $\mathcal{I}$ is given by $\{\bm{y}_{t}\}_{t=1}^{T}$, to compute hypothetical scenarios in-sample for $\tau\in\{1,2,\ldots,T\}$ using the distribution of $\bm{y}_{\tau+h}\>|\>\{\bm{y}_{t}\}_{t=1}^{T}$ conditional on parameters informed by the full information set. In either case, the distribution $p(\bm{y}_{\tau+1}\>|\>\mathcal{I})$ in general does not take a well-known form, and neither does the distribution of higher-order forecasts for $h\geq 2$.
However, we may still explore and obtain random samples from them via predictive simulation. This involves exploiting the fact that even though $p(\bm{y}_{\tau+1}\>|\>\mathcal{I})$ is unknown, $p(\bm{y}_{\tau+1}\>|\>\mathcal{I},\bm{\Xi})$ takes a conditionally Gaussian form under the quite flexible model of Equation (1). Let $m$ denote the current iteration of the MCMC algorithm and $x^{(m)}$ indicates the $m$th draw of a random variable. The one-step-ahead predictive distribution is given by:
where $\bm{x}_{\tau+1}=(\bm{y}_{\tau}^{\prime},\ldots,\bm{y}_{\tau-p+1}^{\prime})^{\prime}$. For $h\geq 2$ we may iterate forward, conditioning recursively on the draws for preceding horizons, by setting the predictors to $\bm{x}_{\tau+h}^{(m)}=(\bm{y}_{\tau+h-1}^{(m)}{{}^{\prime}},\bm{y}_{\tau+h-2}^{(m)}{{}^{\prime}},\ldots)^{\prime}$, and obtain:
where $\bm{y}_{\tau+1:\tau+h-1}^{(m)}$ denotes the path of the variables from $\tau+1$ to $\tau+h-1$ and $\bm{y}_{\tau+1:\tau+h}=(\bm{y}_{\tau+1}^{\prime},\ldots,\bm{y}_{\tau+h}^{\prime})^{\prime}$. This exploits the fact that:
$p(\bm{y}_{\tau+1:\tau+h}\>|\>\mathcal{I})=\int p(\bm{y}_{\tau+1}\>|\>\mathcal{I},\bm{\Xi})\prod_{j=2}^{h}p(\bm{y}_{\tau+j}\>|\>\bm{y}_{\tau+1},\ldots,\bm{y}_{\tau+j-1},\mathcal{I},\bm{\Xi})p(\bm{\Xi}\>|\>\mathcal{I})d\bm{\Xi},$ (6)
that is, the joint distribution of forecasts can be decomposed into the product of the conditional one-step ahead predictive densities. Simulating the process forward, by sampling from the distribution in Equation (5) across horizons $h=1,2,\ldots,$ in each sweep of our algorithm, delivers draws from $p(\bm{y}_{\tau+1:\tau+h}\>|\>\mathcal{I})$, via Monte Carlo integration.
3.1. Conditional forecasts
Suppose we want to impose restrictions on a predefined path of one or more variables. Formally, this implies an additional conditioning argument for the predictive distribution, see also Crump et al. (2025) for related discussions. In this context, we denote by $\mathcal{C}_{h}$ a set that defines the desired restrictions at horizon $h=1,2,\ldots,$ i.e., the unconditional forecast results when $\mathcal{C}_{h}=\varnothing$ for all $h$, and $\mathcal{C}_{1:h}=\{\mathcal{C}_{1},\ldots,\mathcal{C}_{h}\}$. In our version, interest centers on the conditional distribution $p(\bm{y}_{\tau+1:\tau+h}\>|\>\mathcal{I},\mathcal{C}_{1:h})$, which in line with Equation (6) can written as:
[…]
$p(\bm{y}_{\tau+1:\tau+h}\>|\>\mathcal{I},\mathcal{C}_{1:h},\bm{\Xi})$ $=p(\bm{y}_{\tau+1}\>|\>\mathcal{I},\mathcal{C}_{1},\bm{\Xi})\prod_{j=2}^{h}p(\bm{y}_{\tau+j}\>|\>\bm{y}_{\tau+1},\ldots,\bm{y}_{\tau+j-1},\mathcal{I},\mathcal{C}_{1:j},\bm{\Xi}).$
[…]
Notably this differs from the “traditional” implementation of conditional forecasts, that impose the conditions as $p(\bm{y}_{t+j}\>|\>\mathcal{C}_{1:h},\bullet)$ instead of $p(\bm{y}_{t+j}\>|\>\mathcal{C}_{1:j},\bullet)$ for $j=1,\ldots,h$. Put simply, our approach “filters” forward in a loose sense of the word by conditioning on the history of restrictions, thereby resulting in a future path of the unrestricted variables that is consistent with the imposed restrictions over the full set of horizons. By contrast, versions in the spirit of Waggoner and Zha (1999) condition on future restrictions as well as past (and current) ones at each horizon, either by drawing the entire paths of the reduced form shocks or by smoothing via backwards recursions (see Bańbura et al., 2015, for an excellent discussion).
In static multivariate problems these two approaches coincide, while, intuitively, they will potentially lead to increasingly different results as the persistence of the underlying dynamic processes increases. We explore this apparent limitation and the practical usefulness of our approach below in an exercise using artificial data. For dynamic systems featuring a small to moderate amount of persistence (e.g., as in typical applications with monthly/quarterly macroeconomic and financial data that are transformed towards approximate stationarity), the two versions’ outputs are usually close.
Constraints. We operationalize two distinct versions of the restrictions. Define the selection matrix $\bm{R}_{h}$ of size $r_{h}\times n$, where $r_{h}$ is the number of restrictions which may vary across horizons. This matrix serves to select the respective restricted variables. There are two main cases of importance: hard and soft constraints. In the former case, some variables exactly follow a predetermined path; in the latter, they lie within a predetermined interval. As this interval narrows, the soft-constrained case approaches the hard-constrained one. We store these restrictions in the vector $\bm{r}_{h}$ of size $r_{h}\times 1$, whereas $\underline{\bm{r}_{h}}$ and $\overline{\bm{r}_{h}}$, both of size $r_{h}\times 1$, denote the lower and upper bounds of an interval. For hard constraints, we have an equality restriction (see Equation 7), while the soft constraints require an inequality restriction (see Equation 8):
$\mathcal{C}_{h}$ =$=\{\bm{R}_{h}\bm{y}_{\tau+h}=\bm{r}_{h}\},$ (7)
To simplify notation, we omit the $m$-superscript that labels MCMC draws in what follows, but stress that the following computations are carried out in each sweep of our sampling algorithm, thereby marginalizing over the parameters. In addition, define the matrix $\bm{U}_{h}$ of dimension $(n-r_{h})\times n$, which mirrors the selection matrix $\bm{R}_{h}$ and is used to select all unrestricted variables. Define the restricted, $\bm{y}_{\tau+h}^{(R)}=\bm{R}_{h}\bm{y}_{\tau+h}$, and unrestricted, $\bm{y}_{\tau+h}^{(U)}=\bm{U}_{h}\bm{y}_{\tau+h}$, subvectors of $\bm{y}_{\tau+h}$. Similar to the derivation in Section 2.1 and as in Chan et al. (2023), using Equation (5), we may then write:
The inequality restrictions of Equation (8) imply that the conditional forecast of interest, $p(\bm{y}_{\tau+h}\>|\>\underline{\bm{r}_{h}}\leq\bm{y}_{\tau+h}^{(R)}\leq\overline{\bm{r}_{h}},\bullet)$, follows a truncated multivariate Gaussian distribution. Since typically only a subset of variables is restricted (i.e., $r_{h}\leq n$), Chan et al. (2025) suggest a blocked updating scheme to unlock computational advantages. Note that one may decompose the joint distribution of the restricted and unrestricted variables into two parts, and first sample from the $r_{h}$-dimensional (rather than $n$-dimensional) truncated multivariate Gaussian:
with moments $\bm{S}^{(R)}_{\tau+h}=\bm{R}_{h}\bm{\Sigma}_{\tau+h}\bm{R}_{h}^{\prime}$ and $\bm{m}^{(R)}_{\tau+h}=\bm{R}_{h}\bm{F}(\bm{x}_{\tau+h})$. An iid draw from this distribution, $\bm{y}_{\tau+h}^{(R,m)}$, can be obtained from this distribution using the minimax tilting method of Botev (2017). On may then set $\bm{r}_{h}=\bm{y}_{\tau+h}^{(R,m)}$, and use Equation (9) to draw the unconstrained forecasts.
Simulation. To assess the performance and implications of our recursive approach relative to alternative “traditional” implementations of conditional forecasts, we provide a comparison with the precision sampler proposed in Chan et al. (2025). Assuming the conditional mean function to be linear allows to contrast both algorithms one-to-one (we cannot use the precision sampler for nonlinear implementations). For this purpose, we simulate artificial data from a linear VAR process with $n=6$, $p=4$ and a maximum horizon $h=20$ for different levels of persistence (which we measure through spectral radius $\rho(\tilde{\bm{A}})$, i.e., the maximum absolute eigenvalues of the companion matrix of the VAR, $\tilde{\bm{A}}$). We label $\rho(\tilde{\bm{A}})\approx 0.6$ as moderate persistence, and $\rho(\tilde{\bm{A}})\approx 0.95$ as high persistence. The scenario hard constraints treats the first three variables as missing and conditions on the realizations of the remaining variables; soft contraint restricts one of the variables to the predefined interval $(\mu_{y}-0.1,\mu_{y}+0.1)$ where $\mu_{y}=1/h\sum_{t=T-h+1}^{T}y_{i,t}$, i.e., the mean of the respective variable over its final $h$ in-sample periods.
(a) hard constraints, moderate persistence
(b) soft constraint, moderate persistence
(c) hard constraints, high persistence
(d) soft constraint, high persistence
The resulting conditional forecasts are shown in Figure 1. The solid grey lines mark the 68 percent credible set and the median of the predictive distribution when using our proposed recursive approach; the dashed navy blue lines show the same percentiles when using the precision sampler. Considering panels (a) and (b), when facing a moderate amount of persistence, the two approaches yield virtually identical estimates. Some deviations are visible in panels (c) and (d), especially for the left-most panel under the hard constraints and for the first two variables under the soft restriction. In the majority of our experiments, the differences were practically negligible. But this is not necessarily the case in all conceivable scenarios.
3.2. Generalized impulse responses
The IRF is a ubiquitous object of interest in macroeconomics and is widely used for both academic and policy analysis. For the purposes of our paper, we follow the literature and define IRFs generically as:
$\frac{\partial\bm{y}_{\tau+h}}{\partial u_{j\tau}}=\mathbb{E}(\bm{y}_{\tau+h}\>|\>\mathcal{I},u_{j\tau}=d)-\mathbb{E}(\bm{y}_{\tau+h}\>|\>\mathcal{I}).$ (10)
As their name suggests, IRFs trace the dynamic evolution of one or more variables in response to an exogenous structural shock impact (the “impulse”) of size $d$ at time $\tau$ across horizons $h$, where the structural shock is encoded in $u_{j\tau}$. Conventional IRFs from (conditionally) linear VARs are symmetric and proportional with respect to the sign and size of shocks, and they are time-invariant, at least in the case of constant parameter VARs. By contrast, in nonlinear models such as those we consider in this paper, the sign, size, and timing of a shock may matter and can lead to nonlinearities in responses. Due to this state dependence, we cannot use standard methods but need to resort to GIRFs instead, see Koop et al. (1996). We discuss our implementation of GIRFs below and then extend the baseline framework to allow for studying various types of counterfactual shock and transmission scenarios.
The model outlined in Equation (1) is conditionally Gaussian, and standard approaches from the SVAR literature (e.g., zero/sign restrictions, instrumental/proxy variables) can be used to identify structural shocks. Suppose we have (either internally or externally) identified the impact of the structural shock $u_{j\tau}$ of size $d=1$ on the endogenous variables of interest, $\partial\bm{y}_{\tau}/\partial u_{j\tau}=\bm{\delta}_{0}^{(m)}$ at iteration $m$. To sample from the posterior distribution of the GIRFs across horizons, we will use the conditionally Gaussian form of Equation (3). Our goal here is to compare factual and counterfactual scenarios, i.e., we want to compare the case when a structural shock occurs with a non-shock baseline scenario. We construct these scenarios by first defining two distinct versions of initial conditions. In particular, $\bm{x}_{\tau+1}^{(m,d)}=((\bm{y}_{\tau}+d\bm{\delta}_{0}^{(m)})^{\prime},\bm{y}_{\tau-1}^{\prime},\ldots,\bm{y}_{\tau-p+1}^{\prime})^{\prime}$, with superscript $d$ indicating the shock scenario for each iteration $m$ of our algorithm, and $\bm{x}_{\tau+1}^{(m,\varnothing)}=(\bm{y}_{\tau}^{\prime},\bm{y}_{\tau-1}^{\prime},\ldots,\bm{y}_{\tau-p+1}^{\prime})^{\prime}$ with superscript $\varnothing$ referring to the “no-shock” scenario, i.e., it uses the actual configuration of the input vector at time $\tau$. Let $\bm{J}$ be a $k\times n$ selection matrix which has an identity matrix in its upper $n\times n$ block and zeroes everywhere else.
The GIRF on impact reflects our desired shock scenario and is given by:
$\bm{\lambda}_{\tau,0}^{(m,d)}=\bm{J}^{\prime}(\bm{x}_{\tau+1}^{(m,d)}-\bm{x}_{\tau+1}^{(m,\varnothing)})=d\bm{\delta}_{0}^{(m)}.$
That is, we assume that the impact of the respective shock is constant over time, and shocks of different sizes can be simulated by setting $d$ accordingly. Using Equation (4), we obtain two predictive distributions of interest:
$p(\bm{y}_{\tau+1}\>|\>\mathcal{I},\bm{\Xi}^{(m)})$ $=\bm{F}^{(m)}(\bm{x}_{\tau+1}^{(m,\varnothing)})+\bm{\epsilon}_{\tau+1},$
where $\bm{\epsilon}_{\tau+1}\sim\mathcal{N}(\bm{0}_{n},\bm{\Sigma}_{\tau+1}^{(m)})$. In line with the definition of an IRF in Equation (10), at horizon $h=1$, we have $\bm{\lambda}_{\tau,1}^{(m,d)}=\bm{F}^{(m)}(\bm{x}_{\tau+1}^{(m,d)})-\bm{F}^{(m)}(\bm{x}_{\tau+1}^{(m,\varnothing)})$. For higher-order responses $h\geq 2$, we may iterate forward, sampling from the distribution of the future (reduced form) shocks to obtain draws from the unconditional predictive distribution and the distribution conditioning on a shock at time $\tau$ for each horizon. Specifically, for $s\in\{d,\varnothing\}$ and analogous to Equation (5), we have $\bm{x}_{\tau+h}^{(m,s)}=(\bm{y}_{\tau+h-1}^{(m,s)}{{}^{\prime}},\bm{y}_{\tau+h-2}^{(m,s)}{{}^{\prime}},\ldots)^{\prime}$ and thus,
which we use to compute the GIRF at generic horizon $h$:
$\bm{\lambda}_{\tau,h}^{(m,d)}$ $=\mathbb{E}(\bm{y}_{\tau+h}\>|\>\mathcal{I},u_{j\tau}=d,\bm{y}_{\tau+1:\tau+h-1}^{(m,d)},\bm{\Xi}^{(m)})-\mathbb{E}(\bm{y}_{\tau+h}\>|\>\mathcal{I},\bm{y}_{\tau+1:\tau+h-1}^{(m,\varnothing)},\bm{\Xi}^{(m)})$ (12)
$=\bm{F}^{(m)}(\bm{x}_{\tau+h}^{(m,d)})-\bm{F}^{(m)}(\bm{x}_{\tau+h}^{(m,\varnothing)}).$
Two main comments about this framework are in order. First, for the case when $\bm{F}$ is a linear function, our approach produces conventional IRFs. Second, when $\bm{F}$ is nonlinear, one cannot apply expectations as straightforwardly as in the linear case. The reason is that when applying the logic of standard recursive substitutions in this context, the presence of nonlinearities complicates the moments of higher-order responses. For instance, taking $h=2$ as an example, when $\bm{F}$ is a linear mapping with associated (companion form) coefficients, $\tilde{\bm{A}}$, we obtain $\mathbb{E}(\tilde{\bm{A}}^{2}\bm{x}_{\tau+1}+\tilde{\bm{A}}\bm{\epsilon}_{\tau+1}+\bm{\epsilon}_{\tau+2})=\tilde{\bm{A}}^{2}\bm{x}_{\tau+1}$. That is, the IRF at horizon $h$ is given by $\bm{J}^{\prime}\tilde{\bm{A}}^{h}(\bm{x}_{\tau+1}^{(m,d)}-\bm{x}_{\tau+1}^{(m,\varnothing)})$ where $\bm{x}_{\tau+1}^{(m,d)}-\bm{x}_{\tau+1}^{(m,\varnothing)}=(d\bm{\delta}_{0}^{(m)}{{}^{\prime}},\bm{0}_{n(p-1)}^{\prime})^{\prime}$, and can be obtained by simply projecting the shock impact forward using powers of $\tilde{\bm{A}}$. Since $d$ can be factored out and the initial conditions cancel, the IRFs are time-invariant, symmetric, and proportional for different shock sizes and signs. By contrast, in a more general and possibly nonlinear context, this is clearly not the case since we obtain $\mathbb{E}[\bm{F}(\bm{F}(\bm{x}_{\tau+1}+\bm{\epsilon}_{\tau+1}))+\bm{\epsilon}_{\tau+2}]\neq\bm{F}(\bm{F}(\bm{x}_{\tau+1}))$, that is, e.g., at $h=2$ the shock dated $\tau+1$ enters nonlinearly, and so on.
Equation (12) thus addresses the reduced form shocks that enter nonlinearly to obtain the respective conditional expectation (implicitly via the recursive simulation of the factual and counterfactual paths of the variables $\bm{y}_{\tau+1:\tau+h-1}^{(m,s)}$). Computing $\bm{\lambda}_{\tau,h}^{(m,d)}$ for each iteration $m$ allows us to explore the posterior of the expression in Equation (10). Notably, we obtain a GIRF across horizons $h=0,1,2,\ldots,$ for each point in time $\tau$. In principle, one may thus consider dynamic time-varying effects of shocks for different subsets of observations (e.g., recessions vs. expansions), or for each period individually. For most of our empirical work, we will focus on simple time averages $\overline{\bm{\lambda}}_{h}^{(m,d)}=1/T\sum_{\tau=1}^{T}\bm{\lambda}_{\tau,h}^{(m,d)}$, and thus abstract from this dimension of asymmetry.
Another possibility for that we explore in our applied work combines our approach to conditional forecasts with our algorithm for computing GIRFs. Specifically, we may switch off specific transmission channels of structural shocks by partially matching the moments of the predictive distribution conditional on the shock with those of the unconditional predictive distribution. By construction, this results in a GIRF that is equal to zero for the restricted dimensions, thereby effectively considering an alternative scenario where the shock cannot propagate through this channel. From an implementation perspective, we may manipulate Equation (11) for $s=d$ using a hard restriction as in Equation (9), such that $\bm{r}_{h}^{(m)}=\bm{R}_{h}\bm{F}^{(m)}(\bm{x}_{\tau+h}^{(m,\varnothing)})$. That is, we impose that $\bm{R}_{h}\bm{\lambda}_{\tau,h}^{(m,d)}=\bm{0}_{r_{h}}$ along the desired $r_{h}$ dimensions.

### v2

@@ 3.1. Multivariate System Estimation
To estimate the multivariate (reduced form) model in Equation (1) efficiently from a computational perspective, we rely on a conditional representation of its $n$ equations. Let $\bm{e}_{i}$ of size $1\times n$ denote the $i$th row of an identity matrix $\bm{I}_{n}$, and $\bm{E}_{i}$ of size $(n-1)\times n$ is the matrix resulting from deleting the $i$th row of $\bm{I}_{n}$. Using $\bm{y}_{-it}=\bm{E}_{i}\bm{y}_{t}$ we may write $\bm{y}_{t}=\bm{e}_{i}^{\prime}y_{it}+\bm{E}_{i}^{\prime}\bm{y}_{-it}$. Under the assumptions of Equation (1) one may derive the conditional distribution: [^6] Related papers often either use a mapping between the structural and reduced form of the VAR to enable equation-by-equation estimation (see, e.g., Hauzenberger et al., 2024c), or rely on factor models for the reduced form errors (see, e.g., Clark et al., 2023). These approaches come with computational and inferential advantages and disadvantages. The former is simple to implement but requires parameterizing a structural form, which may cause issues such as inadvertently (instead of purposefully to achieve structural identification) breaking order-invariance of the equations. The latter allows for order-invariant inference but gives rise to the usual identification challenges of factor models. Our proposed approach uses a reduced form model, and we obtain order-invariant inference. See also Arias et al. (2023); Chan et al. (2024) for related discussions.
@@
which is a Gaussian with variance $\varsigma_{it}^{2}=(\bm{e}_{i}\bm{\Sigma}_{t}^{-1}\bm{e}_{i}^{\prime})^{-1}$ and mean ${\mu}_{it}=\varsigma_{it}^{2}(\bm{e}_{i}\bm{\Sigma}_{t}^{-1}(\bm{F}(\bm{x}_{t})-\bm{E}_{i}^{\prime}\bm{y}_{-it}))$. Indeed, this distribution is equivalent to a more common representation of the conditional multivariate Gaussian (see, e.g., Cong et al., 2017, Section 2). The mean can alternatively be written as ${\mu}_{it}=f_{i}(\bm{x}_{t})+\varsigma_{it}^{2}(\bm{e}_{i}\bm{\Sigma}_{t}^{-1}\bm{E}_{i}^{\prime})(\bm{y}_{-it}-\bm{E}_{i}\bm{F}(\bm{x}_{t}))$, and we define $\tilde{\mu}_{it}=\varsigma_{it}^{2}(\bm{e}_{i}\bm{\Sigma}_{t}^{-1}\bm{E}_{i}^{\prime})(\bm{y}_{-it}-\bm{E}_{i}\bm{F}(\bm{x}_{t}))$, i.e., $\mu_{it}=f_{i}(\bm{x}_{t})+\tilde{\mu}_{it}$.
@@
which can be used in a Gibbs sampler to update the conditional mean relationships by looping through equations $i=1,\ldots,n$. This approach is similar to the one of Esser et al. (2024) and allows to treat each equation of the multivariate system individually, conditional on all other equations. Specifics about how $f_{i}(\bm{x}_{t})$ is estimated, which we turn to next, are virtually irrelevant for using this equation-by-equation estimation algorithm.
@@ 3.2. Bayesian Additive Regression Trees
The approach we discuss in Section 2 works with any implementation of multivariate models with jointly Gaussian errors. That is, assuming a linear functional form for $\bm{F}(\bm{x}_{t})$ combined with suitable priors results in a standard BVAR. [^7] When we consider linear versions of our model for comparisons, we implement this setting with $\bm{F}(\bm{x}_{t})=\bm{A}\bm{x}_{t}$ where $\bm{A}$ is an $n\times k$ matrix of reduced form VAR coefficients. We assume a horseshoe prior with a single global shrinkage parameter on these coefficients, see also Hauzenberger et al. (2024b). Equation (13) can be used to update the VAR coefficients equation-by-equation from their usual Gaussian posteriors. In case we treat $\bm{F}(\bm{x}_{t})$ nonparametrically, several options are available. Due to its versatility and the established favorable empirical properties we mentioned earlier, we use BART to approximate the equation-specific functions in our applied work. That is, we consider a sum of $s=1,\ldots,S,$ tree functions $\ell_{is}(\bm{x}_{t}\>|\>\mathcal{T}_{is},\bm{\mathrm{m}}_{is})$:
@@ 3.4. Posterior Distributions and Sampling Algorithm
on a $t$-by-$t$ basis. The posterior distribution of the outlier probability is $\mathfrak{p}\>|\>\bullet\sim\mathcal{B}(a_{\mathfrak{p}}+T_{\text{o}},b_{\mathfrak{p}}+T-T_{\text{o}})$, with the total number of observations classified as outliers denoted by $T_{\text{o}}=\sum_{t=1}^{T}\mathbb{I}(s_{t}\neq 1)$, where $\mathbb{I}(\bullet)$ is an indicator function that yields $1$ if its argument is true and 00 otherwise. This completes our modeling framework and algorithmic implementation.